# Sheffer polynomials and the s-ordering of exponential boson operators

### v1

@@ 2.2 Orderings
Equation (17) can be written in terms of a two-variable Hermite polynomial [27, 28], as was recognized early (when $s\in\{0,\pm 1\}$) by Agarwal and Wolf [24, Table V]. It can also be written as
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ $=\sum_{k=0}^{\min(m,n)}\frac{1}{k!}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{\left(\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right)^{k}{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}$ (20)
@@
$\mathopen{:}\,\mathrel{F}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F}\,\mathclose{:}_{s^{\prime}}$ (21)
@@
Suppose now that $F^{(s)},F^{(s^{\prime})}\in{\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$ are the $s$-ordered and $s^{\prime}$-ordered “c-number” versions of some $G\in\mathcal{W}$, computed from $G$ by the requirement (14). By (20), they must be related to each other by the dual relation
$F^{(s)}(\alpha,\alpha^{*})=\exp\left[\left(\frac{s^{\prime}-s}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F^{(s^{\prime})}(\alpha,\alpha^{*}),$ (22)
in which the exponentiated second-order differential operator has been inverted. In the formally infinite series in (22) obtained by expanding the exponential in a power series, only a finite number of terms will be nonzero. Equivalently, $F=F^{(s)}(\alpha,\alpha^{*})$ satisfies the diffusion or heat equation
$\frac{\partial F}{\partial(-s)}=\frac{1}{2}\,\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}F,$ (23)
an unexponentiated version of (22). (See [2, Chapter 12].)
Equation (22) can be extended to the case when $F^{(s)},F^{(s^{\prime})}$ are convergent power series in $\alpha,\alpha^{*}$ rather than polynomials, at least formally. The nicest case is when each of $F^{(s)},F^{(s^{\prime})}$, regarded as a Maclaurin expansion at $(\alpha,\alpha^{*})=(0,0)$, i.e., $(\alpha_{r},\alpha_{i})=(0,0)$, converges globally and defines an analytic function of the real variables $\alpha_{r},\alpha_{i}$, or even a complex analytic function of $\alpha=\alpha_{r}+\nobreak{\rm i}\alpha_{i}$. Because $\partial^{2}/\partial\alpha\partial\alpha^{*}$ equals $\Delta/4$, where $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ is the Laplacian on the complex $\alpha$-plane coordinatized by $\alpha_{r},\alpha_{i}$, one has
$F^{(s)}(\alpha,\alpha^{*})$ $={\rm e}^{(s^{\prime}-s)\Delta/8}F^{(s^{\prime})}(\alpha,\alpha^{*})$ (24)
@@
So in this case, when $s<s^{\prime}$ it may be possible to compute $F^{(s)}$ from $F^{(s^{\prime})}$ by a smoothing Weierstrass transform operation on the complex $\alpha$-plane: convolution with a Gaussian kernel of variance $(s^{\prime}-\nobreak s)/4$.
This operation is familiar from quantum optics [2, 3, 4, 20]: among the quasiprobability densities $F^{(s)}$ obtained when $G=G(a,{a^{\dagger}})$ is a density operator $\hat{\rho}$, the Wigner $W$ function ($s=\nobreak 0$) is a smoothed version of the Glauber–Sudarshan $P$ function ($s=\nobreak+1$), and the Husimi $Q$ function ($s=\nobreak-1$) is a smoothed version of the $W$ function. The smoothing of $P$ to $W$ and $Q$ has been extended from functions of $\alpha\in\mathbb{C}$ to tempered and other distributions on the complex $\alpha$-plane, and the variation with $s$ of the mathematical properties and physical interpretation of smoothed quasiprobability densities has been explored [2, 6, 20].
In the proof of the main result of this paper (Theorem 5.3), the unexponentiated form (23), and not the exponentiated forms (22),(24), will be employed to relate $F^{(s)}$, $F^{(s^{\prime})}\in\mathcal{W}_{c}[[\lambda]]$, to obviate consideration of series convergence.
@@ 3.1 Exponential Riordan arrays and polynomials
Let $d,h\in\mathbb{C}[[z]]$ be formal power series in $z$ of order 0 and $1$ respectively, meaning that their first nonzero terms are the $z^{0}$ and $z^{1}$ terms; so $1/d$ and the compositional inverse $\bar{h}$ both exist as formal power series. (For instance, $d,h$ may be Maclaurin expansions of functions $d(z),h(z)$ analytic at $z=0$.) One usually requires that $d$ be monic, $d=1\cdot z^{0}+\nobreak O(z^{1})$, and often that $h=1\cdot z^{1}+\nobreak O(z^{2})$. Then, a polynomial sequence $s_{n}(t)$, $n\geqslant 0$, where $\deg s_{n}=n$, with EGF $\mathcal{S}(t,z)=d(z){\rm e}^{th(z)}$ that is an element of $\mathbb{C}[t][[z]]$, is defined by
$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (25)
with $s_{0}(t)\equiv 1$. The sequence $s_{n}(t)$ is said to be exponential Riordan for the ordered pair $[d,h]$, and is denoted briefly by $\mathbf{R}[d,h]$. Such notations as $s_{n}(t)=\mathbf{R}[d,h]_{n}(t)$ and $\mathcal{S}(t,z)=\mathbf{R}[d,h](t,z)$ are also used.
@@
$\sum_{n=0}^{\infty}\sum_{k=0}^{n}s_{n,k}t^{k}\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (26)
@@
$s_{n,k}=\frac{n!}{k!}\,[z^{n}]\left[d(z)h(z)^{k}\right],$ (27)
where $z^{n}]$ extracts the coefficient of $z^{n}$ from the element of $\mathbb{C}[[z]]$ that follows. Then $s_{n}(t)=\sum_{k=0}^{n}s_{n,k}t^{k}$, so the $s_{n}(t)$ are the row polynomials of the array. The array $s_{n,k}$ can be viewed as the lower-triangular part of an infinite matrix, and is called an exponential Riordan array. The notation $s_{n,k}=\mathbf{R}[d,h]_{n,k}$ is also used, with $\mathbf{R}[d,h]$ alternatively signifying the matrix as a whole.
@@
In the exponential Riordan group $\mathfrak{R}$ the product operation is $\mathbf{R}[d_{1},h_{1}]\,\mathbf{R}[d_{2},h_{2}]=\allowbreak\mathbf{R}[(d_{2}\circ\nobreak h_{1})d_{1},\allowbreak h_{2}\circ\nobreak h_{1}]$, and the inversion operation is $\mathbf{R}[d,h]^{-1}=\allowbreak\mathbf{R}[1/(d\circ\nobreak\bar{h}),\bar{h}]$, where $\bar{h}$ is the compositional inverse of $h$, i.e., $\bar{h}(h(z))=\nobreak z$. The multiplicative identity is $[1,z]$, which signifies the infinite identity matrix.
@@
$\mathbf{R}[d,h]^{-1}\mathrel{=:}\mathbf{R}\overline{[d,h]}=\mathbf{R}[\bar{d},\bar{h}],$ (28)
@@
The generating functions of the theory facilitate the computation of matrix–vector products.
@@ 3.2 Sheffer polynomials
$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\bar{g}(z){\rm e}^{t\bar{f}(z)}=\frac{1}{g(\bar{f}(z))}{\rm e}^{t\bar{f}(z)},$ (30)
@@
Many well-known polynomial sequences can be viewed as Sheffer sequences [9, 38], and there is some overlap with the orthogonal polynomials of theoretical physics. An example is the sequence of Hermite polynomials $\mathop{{}\it He}\nolimits_{n}(t)$, which is $\mathbf{S}[{\rm e}^{D^{2}/2},D]$ or equivalently $\mathbf{R}[{\rm e}^{-z^{2}/2},z]$. The $\mathop{{}\it He}\nolimits_{n}(t)$ are orthogonal with respect to an appropriate measure (namely, ${\rm e}^{-t^{2}/2}\,{\rm d}t$), as they happen to satisfy a three-term recurrence relation. The squared Hermite polynomials $\mathop{{}\it He}\nolimits_{n}(\sqrt{t})]^{2}$ turn out to be $\mathbf{S}[\sqrt{1-\nobreak 2D}/\allowbreak(1-\nobreak D),\allowbreak D/(1-\nobreak D)]$ or $\mathbf{R}[1/\sqrt{1-\nobreak z^{2}},\allowbreak z/\allowbreak(1+\nobreak z)]$, but do not satisfy such a relation.
For a Sheffer sequence $\mathbf{S}[g,f]$, the formal power series $z=f(D)$ may be the Maclaurin expansion of an elementary function of $D$, without the compositional inverse $D=\bar{f}(z)$ appearing in the EGF (30) having the same property. An example is the sequence of Abel polynomials $A_{n}(t)=t(t-\nobreak n)^{n-1}$, which turns out to be ${\bf S}[1,D{\rm e}^{D}]$. (See [9].) But the compositional inverse of $z=f(D)=D{\rm e}^{D}$ is the Lambert $W$ function $D=\bar{f}(z)=\allowbreak W_{L}(z)=\allowbreak z-\nobreak z^{2}+\nobreak\frac{3}{2}z^{3}+\nobreak\cdots$, which is not an elementary function. It is more convenient to refer to the Abel sequence as ${\bf S}[1,D{\rm e}^{D}]$ than as ${\bf R}[1,W_{L}(z)]$. The same phenomenon will appear in Sec. 5.
@@
$P=g(D),\qquad M=\left[t-\frac{g^{\prime}(D)}{g(D)}\right]\frac{1}{f^{\prime}(D)},$
@@
an equality in $\mathbb{C}[t][[\lambda]]$ which like (31) can be used as an alternative definition of the polynomials $s_{n}$ [9].
The exponentiated raising operator ${\rm e}^{\lambda M}$ is of interest. Each non-negative power $M^{n}$ is an element of $\mathbb{C}[t][[D_{t}]]$, and because $M^{n}s_{0}=s_{n}$ and $s_{0}(t)\equiv 1$, it follows from (30) that
$\left[{\rm e}^{\lambda M}1\right](t)=\bar{g}(\lambda){\rm e}^{t\bar{f}(\lambda)},$ (32)
@@ 3.3 Hsu–Shiue arrays and polynomials
$x^{n}=\sum_{k=0}^{n}\left\{{n\atop k}\right\}x^{\underline{k}},$ (33)
where $x^{\underline{k}}=x(x-1)\dots(x-k+1)$ is the falling factorial with unit “stride.” So, they are coefficients of connection between bases of $\mathbb{C}[x]$. The sequence of row polynomials $\sum_{k=0}^{n}\left\{{n\atop k}\right\}t^{k}$, $n\geqslant 0$, which could be called Stirling polynomials (of the second kind), but are more often called Touchard, Bell, or exponential polynomials, is the exponential Riordan sequence ${\bf R}[1,\allowbreak{\rm e}^{z}-\nobreak 1]$ or equivalently the Sheffer sequence ${\bf S}[1,\allowbreak\log(1+\nobreak D)]$.
@@
$(x+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,x^{\underline{k},B},$ (34)
where $(x+r)^{\underline{n},A}$, $x^{\underline{k},B}$ are falling factorials with stride not necessarily equal to unity, e.g., $x^{\underline{k},B}=x(x-\nobreak B)\dots[x-\nobreak(k-\nobreak 1)B]$. The $\mathop{{}\it HS}\nolimits_{n,k}$ reduce to $\left\{{n\atop k}\right\}$ if $(A,B,r)=(0,1,0)$. The parametric row polynomials
$\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)t^{k},\qquad n\geqslant 0,$ (35)
@@
the right-hand sides being elements of $\mathbb{C}[[z]]$. (The cases $A=0$, $B=0$ are handled by taking limits.) One way of confirming this is to verify that eq. (34), viewed as an equality between a column vector indexed by $n$ and a matrix–vector product, is consistent with Proposition 3.2.
@@
$\overline{[d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]}={[d_{\textit{HS}}(\cdot;B,-r),h_{\textit{HS}}(\cdot;B,A)]},$ (37)
@@
$\begin{aligned} \mathcal{HS}(A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]\end{aligned}$ (38)
@@
$\begin{aligned} \mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]_{n}(t)\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]_{n}(t).\end{aligned}$ (39)
@@
$\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)=\frac{1}{B^{k}k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}(Bj+r)^{\underline{n},A},$ (40)
which comes from eq. (34) by finite-differencing to extract coefficients [40]. And for any $A,B,r$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can be computed from the recurrence (42), below.
@@
    $\left[(-Az-1)\frac{\partial}{\partial z}+Bt\,\frac{\partial}{\partial t}+(r+t)\right]\mathcal{HS}=0,$ (41)
@@
    $\mathop{{}\it HS}\nolimits_{n+1,k+1}=\left[-An+B(k+1)+r\right]\mathop{{}\it HS}\nolimits_{n,k+1}+\mathop{{}\it HS}\nolimits_{n,k},$ (42)
@@
That $\mathcal{HS}(t,z)$ as supplied by Proposition 3.4 satisfies the PDE follows by direct computation. The PDE (with initial condition) can be solved on a neighborhood of $(t,z)=(0,0)$ by the Lagrange–Charpit method of characteristics, yielding the same analytic formula for $\mathcal{HS}(t,z)$. And, it is easily checked that $\mathcal{HS}(t,z)$, a function analytic at $(0,0)$ with bivariate Maclaurin expansion in $\mathbb{C}[t][[z]]$, satisfies the PDE if and only if its expansion coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ satisfy the recurrence (42). ∎
For some choices of $A,B,r$, explicit expressions for the polynomial coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ and the polynomials $\mathop{{}\it HS}\nolimits_{n}(t;A,B,r)$ can be obtained from the summation formula (40) or the recurrence (42). For instance,
$\mathop{{}\it HS}\nolimits_{n,k}(B,B,r)=\binom{n}{k}r^{\underline{n-k},B},$ (43a)
$\mathop{{}\it HS}\nolimits_{n,k}(-B,B,r)=\binom{n}{k}(Bk+r)^{\overline{n-k},B},$ (43b)
with an obvious notation for the rising factorial. Notice that by (43b),
$\mathop{{}\it HS}\nolimits_{n,k}(-1,1,1)={\binom{n}{k}}\frac{n!}{k!},$ $\mathop{{}\it HS}\nolimits_{n,k}(1,-1,-1)=(-1)^{n-k}{\binom{n}{k}}\frac{n!}{k!},$ (44)

### v2

@@ 2.2 Orderings
Equation (2.13) can be written in terms of a two-variable Hermite polynomial [27, 28], as was recognized early (when $s\in\{0,\pm 1\}$) by Agarwal and Wolf [24, Table V]. It can also be written as
@@
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ $=\sum_{k=0}^{\min(m,n)}\frac{1}{k!}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{\left(\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right)^{k}{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}$ (2.16)
@@
$\mathopen{:}\,\mathrel{F}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F}\,\mathclose{:}_{s^{\prime}}$ (2.17)
@@
Suppose now that $F^{(s)},F^{(s^{\prime})}\in{\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$ are the $s$-ordered and $s^{\prime}$-ordered “c-number” versions of some $G\in\mathcal{W}$, computed from $G$ by the requirement (2.10). By (2.16), they must be related to each other by the dual relation
$F^{(s)}(\alpha,\alpha^{*})=\exp\left[\left(\frac{s^{\prime}-s}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F^{(s^{\prime})}(\alpha,\alpha^{*}),$ (2.18)
in which the exponentiated second-order differential operator has been inverted. In the formally infinite series in (2.18) obtained by expanding the exponential in a power series, only a finite number of terms will be nonzero. Equivalently, $F=F^{(s)}(\alpha,\alpha^{*})$ satisfies the diffusion or heat equation
$\frac{\partial F}{\partial(-s)}=\frac{1}{2}\,\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}F,$ (2.19)
an unexponentiated version of (2.18). (See [2, Chapter 12].)
Equation (2.18) can be extended to the case when $F^{(s)},F^{(s^{\prime})}$ are convergent power series in $\alpha,\alpha^{*}$ rather than polynomials, at least formally. The nicest case is when each of $F^{(s)},F^{(s^{\prime})}$, regarded as a Maclaurin expansion at $(\alpha,\alpha^{*})=(0,0)$, i.e., $(\alpha_{r},\alpha_{i})=(0,0)$, converges globally and defines an analytic function of the real variables $\alpha_{r},\alpha_{i}$, or even a complex analytic function of $\alpha=\alpha_{r}+\penalty 10000{\rm i}\alpha_{i}$. Because $\partial^{2}/\partial\alpha\partial\alpha^{*}$ equals $\Delta/4$, where $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ is the Laplacian on the complex $\alpha$-plane coordinatized by $\alpha_{r},\alpha_{i}$, one has
$F^{(s)}(\alpha,\alpha^{*})$ $={\rm e}^{(s^{\prime}-s)\Delta/8}F^{(s^{\prime})}(\alpha,\alpha^{*})$ (2.20)
@@
So in this case, when $s<s^{\prime}$ it may be possible to compute $F^{(s)}$ from $F^{(s^{\prime})}$ by a smoothing Weierstrass transform operation on the complex $\alpha$-plane: convolution with a Gaussian kernel of variance $(s^{\prime}-\penalty 10000s)/4$.
[…]
In the proof of the main result of this paper (Theorem 5.3), the unexponentiated form (2.19), and not the exponentiated forms (2.18),(2.20), will be employed to relate $F^{(s)}$, $F^{(s^{\prime})}\in\mathcal{W}_{c}[[\lambda]]$, to obviate consideration of series convergence.
@@ 3.1 Exponential Riordan arrays and polynomials
Let $d,h\in\mathbb{C}[[z]]$ be formal power series in $z$ of order 0 and $1$ respectively, meaning that their first nonzero terms are the $z^{0}$ and $z^{1}$ terms; so $1/d$ and the compositional inverse $\bar{h}$ both exist as formal power series. (For instance, $d,h$ may be Maclaurin expansions of functions $d(z),h(z)$ analytic at $z=0$.) One usually requires that $d$ be monic, $d=1\cdot z^{0}+\penalty 10000O(z^{1})$, and often that $h=1\cdot z^{1}+\penalty 10000O(z^{2})$. Then, a polynomial sequence $s_{n}(t)$, $n\geqslant 0$, where $\deg s_{n}=n$, with EGF $\mathcal{S}(t,z)=d(z){\rm e}^{th(z)}$ that is an element of $\mathbb{C}[t][[z]]$, is defined by
$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (3.1)
with $s_{0}(t)\equiv 1$. The sequence $s_{n}(t)$ is said to be exponential Riordan for the ordered pair $[d,h]$, and will be denoted briefly by $\mathbf{R}[d,h]$. Such notations as $s_{n}(t)=\mathbf{R}[d,h]_{n}(t)$ and $\mathcal{S}(t,z)=\mathbf{R}[d,h](t,z)$ are also used.
@@
$\sum_{n=0}^{\infty}\sum_{k=0}^{n}s_{n,k}t^{k}\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (3.2)
@@
$s_{n,k}=\frac{n!}{k!}\,[z^{n}]\left[d(z)h(z)^{k}\right],$ (3.3)
where $z^{n}]$ extracts the coefficient of $z^{n}$ from the element of $\mathbb{C}[[z]]$ that follows. Then $s_{n}(t)=\sum_{k=0}^{n}s_{n,k}t^{k}$, so the $s_{n}(t)$ are the row polynomials of the array. The array $s_{n,k}$ can be viewed as the lower-triangular part of an infinite matrix, and is called an exponential Riordan array. The notation $s_{n,k}=\mathbf{R}[d,h]_{n,k}$ will also be used, with $\mathbf{R}[d,h]$ alternatively signifying the matrix as a whole.
@@
In the exponential Riordan group $\mathfrak{R}$ the product operation is $\mathbf{R}[d_{1},h_{1}]\,\mathbf{R}[d_{2},h_{2}]=\allowbreak\mathbf{R}[(d_{2}\circ\penalty 10000h_{1})d_{1},\allowbreak h_{2}\circ\penalty 10000h_{1}]$, and the inversion operation is $\mathbf{R}[d,h]^{-1}=\allowbreak\mathbf{R}[1/(d\circ\penalty 10000\bar{h}),\bar{h}]$, where $\bar{h}$ is the compositional inverse of $h$, i.e., $\bar{h}(h(z))=\penalty 10000z$. The multiplicative identity is $[1,z]$, which signifies the infinite identity matrix.
@@
$\mathbf{R}[d,h]^{-1}\mathrel{=:}\mathbf{R}\overline{[d,h]}=\mathbf{R}[\bar{d},\bar{h}],$ (3.4)
@@
A quantum-optics example of inversion in $\mathfrak{R}$ is afforded by the Cahill–Glauber identity (1.2) and its inverse (1.3): expansion in the parameter $\lambda$ reveals that they are based respectively on the arrays
$\mathbf{R}\left[\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}\right],\quad\mathbf{R}\left[\frac{2}{2-z-sz},\ln\left(\frac{2+z-sz}{2-z-sz}\right)\right]$ (3.5)
in $\mathfrak{R}$, which are inverses of each other.
The generating functions of the Riordan-array theory facilitate the computation of matrix–vector products, as follows.
@@ 3.2 Sheffer polynomials
$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\bar{g}(z){\rm e}^{t\bar{f}(z)}=\frac{1}{g(\bar{f}(z))}{\rm e}^{t\bar{f}(z)},$ (3.7)
@@
[…]
For a Sheffer sequence $\mathbf{S}[g,f]$, the formal power series $z=f(D)$ may be the Maclaurin expansion of an elementary function of $D$, without the compositional inverse $D=\bar{f}(z)$ appearing in the EGF (3.7) having the same property. An example is the sequence of Abel polynomials $A_{n}(t)=t(t-\penalty 10000n)^{n-1}$, which turns out to be ${\bf S}[1,D{\rm e}^{D}]$. (See [9].) But the compositional inverse of $z=f(D)=D{\rm e}^{D}$ is the Lambert $W$ function $D=\bar{f}(z)=\allowbreak W_{L}(z)=\allowbreak z-\penalty 10000z^{2}+\penalty 10000\frac{3}{2}z^{3}+\penalty 10000\cdots$, which is not an elementary function. It is more convenient to refer to the Abel sequence as ${\bf S}[1,D{\rm e}^{D}]$ than as ${\bf R}[1,W_{L}(z)]$. The same phenomenon will appear in Sec. 5.
@@
$P=f(D),\qquad M=\left[t-\frac{g^{\prime}(D)}{g(D)}\right]\frac{1}{f^{\prime}(D)},$
@@
an equality in $\mathbb{C}[t][[\lambda]]$ which like (3.8) can be used as an alternative definition of the polynomials $s_{n}$ [9].
The exponentiated raising operator ${\rm e}^{\lambda M}$ is of interest. Each non-negative power $M^{n}$ is an element of $\mathbb{C}[t][[D_{t}]]$, and because $M^{n}s_{0}=s_{n}$ and $s_{0}(t)\equiv 1$, it follows from (3.7) that
$\left[{\rm e}^{\lambda M}1\right](t)=\bar{g}(\lambda){\rm e}^{t\bar{f}(\lambda)},$ (3.9)
@@ 3.3 Hsu–Shiue arrays and polynomials
$x^{n}=\sum_{k=0}^{n}\left\{{n\atop k}\right\}x^{\underline{k}},$ (3.10)
[…]
@@
$(x+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,x^{\underline{k},B},$ (3.11)
[…]
$\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)t^{k},\qquad n\geqslant 0,$ (3.12)
@@
the right-hand sides being elements of $\mathbb{C}[[z]]$. (The cases $A=0$, $B=0$ are handled by taking limits.) One way of confirming this is to verify that eq. (3.11), viewed as an equality between a column vector indexed by $n$ and a matrix–vector product, is consistent with Proposition 3.2.
@@
$\overline{[d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]}={[d_{\textit{HS}}(\cdot;B,-r),h_{\textit{HS}}(\cdot;B,A)]},$ (3.14)
@@
$\begin{aligned} \mathcal{HS}(A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]\end{aligned}$ (3.15)
@@
$\begin{aligned} \mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]_{n}(t)\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]_{n}(t).\end{aligned}$ (3.16)
@@
$\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)=\frac{1}{B^{k}k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}(Bj+r)^{\underline{n},A},$ (3.17)
which comes from eq. (3.11) by finite-differencing to extract coefficients [40]. And for any $A,B,r$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can be computed from the recurrence (3.19), below.
@@
    $\left[(-Az-1)\frac{\partial}{\partial z}+Bt\,\frac{\partial}{\partial t}+(r+t)\right]\mathcal{HS}=0,$ (3.18)
@@
    $\mathop{{}\it HS}\nolimits_{n+1,k+1}=\left[-An+B(k+1)+r\right]\mathop{{}\it HS}\nolimits_{n,k+1}+\mathop{{}\it HS}\nolimits_{n,k},$ (3.19)
@@
That $\mathcal{HS}(t,z)$ as supplied by Proposition 3.4 satisfies the PDE follows by direct computation. The PDE (with initial condition) can be solved on a neighborhood of $(t,z)=(0,0)$ by the Lagrange–Charpit method of characteristics, yielding the same analytic formula for $\mathcal{HS}(t,z)$. And, it is easily checked that $\mathcal{HS}(t,z)$, a function analytic at $(0,0)$ with bivariate Maclaurin expansion in $\mathbb{C}[t][[z]]$, satisfies the PDE if and only if its expansion coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ satisfy the recurrence (3.19). ∎
For some choices of $A,B,r$, explicit expressions for the polynomial coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ and the polynomials $\mathop{{}\it HS}\nolimits_{n}(t;A,B,r)$ can be obtained from the summation formula (3.17) or the recurrence (3.19). For instance,
$\mathop{{}\it HS}\nolimits_{n,k}(B,B,r)=\binom{n}{k}r^{\underline{n-k},B},$ (3.20a)
$\mathop{{}\it HS}\nolimits_{n,k}(-B,B,r)=\binom{n}{k}(Bk+r)^{\overline{n-k},B},$ (3.20b)
with an obvious notation for the rising factorial. Notice that by (3.20b),
$\mathop{{}\it HS}\nolimits_{n,k}(-1,1,1)={\binom{n}{k}}\frac{n!}{k!},$ $\mathop{{}\it HS}\nolimits_{n,k}(1,-1,-1)=(-1)^{n-k}{\binom{n}{k}}\frac{n!}{k!},$ (3.21)